The bird in fig A drops the stone from a height of 88 m above the
water surface. Initially, the stone has zero vertical velocity. How
long will it take the stone to reach the surface of the pond?
Assume air resistance is negligible.
1
Expert's answer
2020-11-03T10:31:40-0500

Solution

The time taken by stone to reach on the surface of pond

is second equation of motion

"S=ut+\frac{gt^2}{2}"

Initial velocity is zero

g=9.8m/s2

So

Time taken by stone is


"t=\sqrt{\frac{2S}{g}}"


"t=\sqrt{\frac{2\times88}{9.8}}"


t=4.29 sec.